AB= 14x + 25
BC= 11x - 31
What is x and AC?

Answer by jim_thompson5910(35256)

AC = AB + BC

AC = (14x + 25) + (11x - 31)

AC = 14x + 25 + 11x - 31

AC = 25x - 6

So we need to know the length of AC to know the value of x.

If AB = BC, then...

14x + 25 = 11x - 31

14x - 11x = -31 - 25

3x = -56

x = -56/3

Now use this x value to find AC

Note: x = -56/3 only if AB = BC